# Highly edge-connected subgraphs

### v1

@@ Abstract
Mader conjectured that an average degree of at least 3k13k-1 is sufficient for the existence of a $(k+1)$-connected subgraph. The following minimum degree version holds: Every graph with minimum degree at least 3k13k-1 has a $(k+1)$-connected subgraph on more than $2k$ vertices. Moreover, for triangle-free graphs, already an average degree of at least $2k$ is sufficient for a $(k+1)$-connected subgraph (which has at least $2(k+1)$ vertices).
@@
It is shown that all of these results are sharp in some sense. The results can be used for the decomposition of graphs into two highly (edge-)connected parts.
@@
Technische Universität Ilmenau, August 11, 2025
On highly (vertex-)connected subgraphs
Mader conjectured in [1] that every graph with average degree 3k13k-1 has a $(k+1)$-connected subgraph. This conjecture turned out to be quite hard, since the best known sufficient linear factor was improved in many iterations, but never reached $3$. We consider another version of the problem by demanding even a minimum degree of 3k13k-1. In exchange for the stronger requirement, the $(k+1)$-connected subgraph is enforced to have a non-trivial size of more than $2k$ vertices. This bonus outcome cannot be expected from an average degree of 3k13k-1, since by a construction of Carmesin in [2], there are graphs of average degree $\big{(}3+\frac{1}{3}\big{)}k-\mathcal{O}(k)$ without a $(k+1)$-connected subgraph on more than $2k$ vertices.
@@
$m_{G}(v)\leavevmode\nobreak\ :=\leavevmode\nobreak\ 0\lor\big{(}(3k-1)-d_{G}(v)\big{)}\land k\leavevmode\nobreak\ \in\{0,\dots,k\}\,.$
We prove by induction on $v(G)>2k$ that
$M(G)\leavevmode\nobreak\ :=\leavevmode\nobreak\ \sum_{v\in V(G)}m_{G}(v)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2k^{2}\,.$
@@
We have $d_{G}(s)\leq v(G)-1\leq 3k-1$, and hence $m_{G}(s)\geq\big{(}(3k-1)-(v(G)-1)\big{)}\land k=3k-v(G)$. It is $m_{G}(a)=m_{G}(b)=k$. Hence
$M(G)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \sum_{s}\big{(}3k-v(G)\big{)}+\sum_{a,b}k\leavevmode\nobreak\ \geq\leavevmode\nobreak\ k(3k-v(G))+(v(G)-k)k\leavevmode\nobreak\ =\leavevmode\nobreak\ 2k^{2}\,.$
@@
We use the induction hypothesis for $A$. We have $d_{G}(s)-d_{A}(s)\leq v(B)-k$. Since the function $x\mapsto 0\lor\big{(}(3k-1)-x\big{)}\land k$ is non-expanding, this yields $m_{A}(s)-m_{G}(s)\leq v(B)-k$. Again, $m_{G}(b)=k$. Hence
$M(G)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{a}m_{A}(a)+\sum_{s}\big{(}m_{A}(a)-(v(B)-k)\big{)}+\sum_{b}k$
$=\leavevmode\nobreak\ M(A)-k(v(B)-k)+(v(B)-k)k\leavevmode\nobreak\ =\leavevmode\nobreak\ M(A)\leavevmode\nobreak\ \geq\leavevmode\nobreak 2k^{2}\,.$
@@
$M(G)\leavevmode\nobreak$ $\geq\leavevmode\nobreak\ \sum_{a}m_{A}(a)+\sum_{s}\big{(}m_{A}(s)+m_{B}(s)-2k\big{)}+\sum_{b}m_{B}(b)$
$=\leavevmode\nobreak\ M(A)+M(B)-k(2k)\leavevmode\nobreak\ \geq\leavevmode\nobreak 2k^{2}+2k^{2}-2k^{2}\leavevmode\nobreak\ =\leavevmode\nobreak 2k^{2}\,.$
@@
Proof.  We present a construction for (i), which we will modify for (ii) at the end.
@@
Let $a,b,c\in\mathbb{N}$ which differ by at most $1$ with a+b+c=k1a+b+c=k-1. Let $V_{0}$ be an anticlique on $k$ vertices and let $V_{1},\dots,V_{l}$ be disjoint cliques disjoint to $V_{0}$ on k1k-1 vertices, where $l=3+2a+b$. We partition each $V_{j}$ for $j\geq 1$ into three cliques $V_{j}=A_{j}\cup B_{j}\cup C_{j}$ with $|A_{j}|=a$, $|B_{j}|=b$ and $|C_{j}|=c$.
Let $w_{1},\dots,w_{2a+b}$ be a list of the vertices in $W:=A_{1}\cup B_{1}\cup A_{2}$. For $j\in\{1,2,3\}$, we add all possible edges between $V_{j}$ and $W_{j}:=V_{0}$. For $j\geq 4$, we add all possible edges between $V_{j}$ and the $k$ vertices in  $W_{j}\leavevmode\nobreak\ :=\leavevmode\nobreak\ A_{j-1}\,\cup\,B_{j-2}\,\cup\,C_{j-3}\,\cup\,\{w_{j-3}\}$. For $j=l$, this is shown in Figure 1.
Figure 1:  The graph $H$. Only the edges from $V_{0}$ and $V_{l}$ are shown.
@@
$|X|\leavevmode\nobreak\ =\leavevmode\nobreak\ c+b+c+(k-1)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 1+a+b+c+(k-1)\leavevmode\nobreak\ =\leavevmode\nobreak\ 2k-1\,.$
@@
Now let $j\geq 4$ and assume that there is a $(k+1)$-connected subgraph $G^{\prime}$ on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices in $G_{j}$. Since the neighborhood $W_{j}$ of $V_{j}$ in $G_{j}$ consists of only $k$ vertices, $G^{\prime}$ is either a subgraph of $G_{j}-V_{j}=G_{j-1}$ or of $G_{j}[V_{j}\cup W_{j}]$. The first case is prevented by the induction hypothesis, so assume the second. Since there are no edges between each two of the sets $A_{j-1}$, $B_{j-2}$ and $C_{j-3}$, and those three are separated after deleting the set $V_{j}\cup\{w_{j-3}\}$ of $k$ vertices, $V(G^{\prime})$ must be completely contained in one of the sets $V_{j}\cup\{w_{j-3}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j-3}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j-3}\}\cup C_{j-3}$. But all of them have size at most $(k-1)+1+\big{\lceil}\frac{k-1}{3}\big{\rceil}\leq\big{\lceil}\frac{4}{3}k\big{\rceil}$, a contradiction.
This proves the claimed properties of $H$. We now inductively construct a series $H_{0},H_{1},\dots$ of graphs without a $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices such that in $H_{j}$ there is a set $X_{j}$ of $(2k-2^{j})\lor 0$ vertices of degree 2k22k-2 and all vertices in $V(H_{j})\setminus X_{j}$ have degree at least 3k33k-3. Then for $j\geq\log_{2}2k$, $H_{j}$ has minimum degree at least 3k33k-3.
Let $H_{0}=H$ and $X_{0}=X$ (possibly replenished to size 2k12k-1), which satisfy the conditions. Now assume we have already constructed $H_{j},X_{j}$ and we want to construct $H_{j+1},X_{j+1}$. Let $Y_{j}\subseteq X_{j}$ with $|Y_{j}|=k\land|X_{j}|$. Let $H^{\prime}_{j}$ be a copy of $H_{j}$ with $V(H_{j})\cap V(H^{\prime}_{j})=Y_{j}$ and $H_{j}[Y_{j}]=H^{\prime}_{j}[Y_{j}]$. We define $H_{j+1}=\big{(}V(H_{j})\cup V(H^{\prime}_{j}),\,E(H_{j})\cup E(H^{\prime}_{j})\big{)}$. Since the subgraphs $H_{j}$ and $H^{\prime}_{j}$ of $H_{j+1}$ intersect only in the vertex set $Y_{j}$ of at most $k$ vertices and contain no $(k+1)$-connected subgraph on more than $\big{\lceil}\frac{4}{3}k\big{\rceil}$ vertices, neither $H_{j+1}$ does. For every $y\in Y_{j}$, we have
$d_{H_{j+1}}(y)\leavevmode\nobreak\ =\leavevmode\nobreak\ d_{H_{j}}(y)+d_{H^{\prime}_{j}}(y)-d_{H_{j}[Y_{j}]}(y)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 2(k-1)+2(k-1)-(k-1)\leavevmode\nobreak\ =\leavevmode\nobreak\ 3(k-1)\,.$
Also the vertices that are not contained in $X_{j}$ or its copy, have degree at least 3k33k-3. The remaining vertices from $X_{j}\setminus Y_{j}$ and its copy form the new set $X_{j+1}$ of size
$|X_{j+1}|\leavevmode\nobreak$ = $=\leavevmode\nobreak 2|X_{j}\setminus Y_{j}|\leavevmode\nobreak\ =\leavevmode\nobreak 2(|X_{j}|-k\land|X_{j}|)\leavevmode\nobreak\ =\leavevmode\nobreak 2(|X_{j}|-k)\lor 0$
= $=\leavevmode\nobreak 2\big{(}(2k-2^{j})\lor 0-k\big{)}\lor 0\leavevmode\nobreak\ =\leavevmode\nobreak 2(2k-2^{j}-k)\lor 0\leavevmode\nobreak\ =\leavevmode\nobreak\ (2k-2^{j+1})\lor 0\,.$
@@
For (ii), we replace the complete graphs on each of the sets $A_{j},B_{j},C_{j}$ for $j\geq 1$ by the disjoint union of two complete graphs of size as equal as possible. By doing so, we loose in $H_{0}=H$ a degree $\frac{1}{6}k+\mathcal{O}(1)$ at every vertex. For the construction of $H_{1}$, we may assume $c\leq a$, and we let $Y_{0}=V_{l}\subseteq X$. Hence, $|X_{1}|$ consists of $C_{l-2}\cup B_{l-1}\cup C_{l-1}$ and its copy, which are $2(2c+b)\leq 2(a+b+c)=2k-2$ vertices. Doing so guarantees $d_{H_{j}[Y_{j}]}(y)\leq k-\frac{1}{6}k+\mathcal{O}(1)$ for every $j\geq 0$, which is needed to show for $y\in Y_{j}$ that
$d_{H_{j+1}}(y)\leavevmode\nobreak$ $=\leavevmode\nobreak\ d_{H_{j}}(y)+d_{H^{\prime}_{j}}(y)-d_{H_{j}[Y_{j}]}(y)$
$\geq\leavevmode\nobreak 2\big{(}2k-\tfrac{1}{6}k-\mathcal{O}(1)\big{)}-\big{(}k-\tfrac{1}{6}k+\mathcal{O}(1)\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \big{(}3-\tfrac{1}{6}\big{)}k-\mathcal{O}(1)\,.$
The proof of the non-existence of a $(k+1)$-connected subgraph first works as for (i): The vertex set of a $(k+1)$-connected subgraph must be completely contained in one of the sets $V_{j}\cup\{w_{j-3}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j-3}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j-3}\}\cup C_{j-3}$, for some $j\geq 4$. Assume $V(G^{\prime})\subseteq V_{j}\cup\{w_{j-3}\}\cup A_{j-1}$. The two components of $G[A_{j-1}]$ can be separated by deleting the other $k$ vertices. Hence $G^{\prime}$ cannot contain vertices from both components. The same holds for $A_{j}\subseteq V_{j}$. Together, $V(G^{\prime})\leq 2\big{\lceil}\frac{a}{2}\big{\rceil}+b+c+1\leq 1+a+b+c+1=k+1$, which is too low for being $(k+1)$-connected. The other two cases work similarly. $\Box$
@@
There is a $C<3$ such that, for all $k\in\mathbb{N}$, every graph of minimum degree at least $Ck$ has a $(k+1)$-connected subgraph.
By Theorem On highly (vertex-)connected subgraphs, a suitable constant must satisfy $C\geq 3-\frac{1}{6}$. The conjecture is posed in a way that allows to assume that $k>k_{0}$ for some sufficiently large $k_{0}$: For $k\leq k_{0}$, we have $\big{(}3-\frac{1}{k_{0}}\big{)}k\,\geq\,\big{(}3-\frac{1}{k}\big{)}k\,=\,3k-1$, so Theorem On highly (vertex-)connected subgraphs suffices.
@@
Every triangle-free graph with average degree at least $2k$ has a $(k+1)$-connected subgraph (on at least $2(k+1)$ vertices).
Proof.  First, we prove the well-known statement that, if $G$ is a triangle-free graph on $2l$ vertices, then $e(G)\leq l^{2}$. This is true for $l=1$. If $l>1$ and $E(G)\neq\emptyset$, consider an edge $vw\in E(G)$. Since $G$ is triangle-free, the neighborhoods of $v$ and $w$ are disjoint and hence $d_{G}(v)+d_{G}(w)\leq v(G)=2l$. This yields $e(G)=e(G-v-w)+d_{G}(v)+d_{G}(w)-1\leq(l-1)^{2}+2l-1=l^{2}$.
@@
Now assume $v(G)>2k$. $G$ itself is not $(k+1)$-connected. Hence it has a $k$-separation $(A,B)$ (see Theorem On highly (vertex-)connected subgraphs). By symmetry, we may assume that $v(A)\geq v(B)$.
@@
$e(G)\leavevmode\nobreak\ =\leavevmode\nobreak\ e(G-b)+d_{G}(b)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ e(G-b)+k\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\big{(}v(G)-1-k\big{)}+k\leavevmode\nobreak\ =\leavevmode\nobreak\ k\big{(}v(G)-k\big{)}\,.$
@@
$e(G)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ e(A)+e(B)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ k\big{(}v(A)-k\big{)}+k\big{(}v(B)-k\big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ k\big{(}v(G)-k\big{)}\,.$
@@
This Theorem is best possible, even in the sense of minimum degree condition: There exists a bipartite graph with minimum degree 2k12k-1 without a $(k+1)$-connected subgraph. One can build such a graph by gluing together copies of the bipartite graph $H_{0}:=K_{k,2k-1}$ similarly to the final step of the proof of Theorem On highly (vertex-)connected subgraphs.
@@
Mader proved in [3] that every simple graph $G$ with $e(G)>k\big{(}v(G)-\frac{k+1}{2}\big{)}$ and $v(G)\geq k$ contains a $(k+1)$-edge-connected subgraph. This condition is sharp for every $k$ and every $v(G)$: Consider the graph $G$ on a clique $A$ on $k$ vertices and an anticlique on $V(G)\setminus A$ with all possible edges between $A$ and $V(G)\setminus A$. Clearly, $G$ has not even a subgraph of minimum degree at least $k+1$. Note that, when $v(G)$ tends to $\infty$, the average degree of $G$ tends to $2k$.
We first show that an average degree of at least $2k$ has some stronger implications. For simple graphs in particular, it implies the existence of a $(k+1)$-edge-connected subgraph on more than $2k$ vertices.
@@
Let $G$ be a (not necessarily simple) graph with e(G)>k(v(G)1)e(G)\,>\,k\,(v(G)-1). Then $G$ contains a subgraph $G^{\prime}$ with
@@
    If $G$ is simple, then $v(G^{\prime})>2k$.
Proof.  Let $G^{\prime}$ be a smallest nonempty induced subgraph of $G$ with $e(G^{\prime})\,>\,k\,(v(G^{\prime})-1)$. Let $\mathcal{P}$ be a partition of $V(G^{\prime})$ in more than one class. Then for all $A\in\mathcal{P}$, $e(G[A])\leq k(|A|-1)$. Hence
$e(\mathcal{P})\leavevmode\nobreak$ $=\leavevmode\nobreak\ e(G^{\prime})\,-\sum_{A\in\mathcal{P}}e(G[A])\leavevmode\nobreak\ >\leavevmode\nobreak\ k(v(G^{\prime})-1)\,-\sum_{A\in\mathcal{P}}k(|A|-1)$
$=\leavevmode\nobreak\ k\Big{(}v(G)-1+|\mathcal{P}|-\sum_{A\in\mathcal{P}}|A|\Big{)}\leavevmode\nobreak\ =\leavevmode\nobreak\ k(|\mathcal{P}|-1)\,.$
In particular, it follows from the case $|\mathcal{P}|=2$, that $G^{\prime}$ is $(k+1)$-edge-connected. Moreover, by the Theorem of Tutte [4] / Nash-Williams [5], $G^{\prime}$ contains $k$ pairwise edge-disjoint spanning trees. If $G$ and hence $G^{\prime}$ is simple, then $G^{\prime}$ has more than $2k$ vertices, since otherwise $e(G^{\prime})\,\leq\,\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)\,\leq\,k\,(v(G^{\prime})-1)$. $\Box$
@@
    Every (2k1)(2k-1)-regular simple graph $G$ without a complete subgraph on $2k$ vertices does not contain a subgraph $G^{\prime}$ with $k$ pairwise edge-disjoint spanning trees, because every subgraph $G^{\prime}$ has less than the necessary $k(v(G^{\prime})-1)$ edges:
@@
    • If $v(G^{\prime})=2k$, then $e(G^{\prime})<\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)=k(v(G^{\prime})-1)$.

### v2

# arXiv:2508.07997v2[math.CO] 02 Mar 2026

@@ Abstract
Mader (1979) conjectured that an average degree of at least 3k13k-1 in a graph is sufficient for the existence of a $(k+1)$-connected subgraph. The following minimum degree version holds: Every graph with minimum degree at least 3k13k-1 has a $(k+1)$-connected subgraph on more than $2k$ vertices. Moreover, for triangle-free graphs, already an average degree of at least $2k$ is sufficient for a $(k+1)$-connected subgraph, which has at least $2(k+1)$ vertices.
@@
It is shown that all of these results are sharp in some sense. The results may be applied to decompose graphs into two highly connected or edge-connected parts.
@@
Technische Universität Ilmenau
On highly connected subgraphs
Mader conjectured in [1] that every graph with average degree 3k13k-1 has a $(k+1)$-connected subgraph. This conjecture turned out to be quite hard, since the best known sufficient linear factor was improved in many iterations, but never reached $3$. We consider another version of the problem by demanding even a minimum degree of 3k13k-1. In exchange for the stronger requirement, the $(k+1)$-connected subgraph is enforced to have a non-trivial size of more than $2k$ vertices. This bonus outcome cannot be expected from an average degree of 3k13k-1, since by a construction of Carmesin in [2], there are graphs of average degree $\big(3+\frac{1}{3}\big)k-\mathcal{O}(1)$ without a $(k+1)$-connected subgraph on more than $2k$ vertices.
@@
$m_{G}(v)\penalty 10000\ :=\penalty 10000\ \min\Big\{\max\big\{(3k-1)-d_{G}(v),\,0\big\},\,k\Big\}\penalty 10000\ \in\{0,\dots,k\}\,.$
We prove by induction on $v(G)>2k$ that the total penalty of $G$ satisfies
$M(G)\penalty 10000\ :=\penalty 10000\ \sum_{v\in V(G)}m_{G}(v)\penalty 10000\ \geq\penalty 10000\ 2k^{2}\,.$
@@
We have $d_{G}(s)\leq v(G)-1$, and hence $m_{G}(s)\geq\min\big\{(3k-1)-(v(G)-1),k\big\}=3k-v(G)$.It is $m_{G}(a)=m_{G}(b)=k$. Hence
$M(G)\penalty 10000\ \geq\penalty 10000\ \sum_{s}\big(3k-v(G)\big)+\sum_{a}k+\sum_{b}k\penalty 10000\ =\penalty 10000\ k(3k-v(G))+(v(G)-k)k\penalty 10000\ =\penalty 10000\ 2k^{2}\,.$
@@
We use the induction hypothesis for $A$. Since the function $x\mapsto\min\big\{\max\{3k-1\,-x,\,0\},\,k\big\}$ is non-expanding, we have $m_{A}(s)-m_{G}(s)\leq\big|d_{A}(s)-d_{G}(s)\big|\leq|V(G)\setminus V(A)|=v(B)-k$. Again, $m_{G}(b)=k$. Hence
[…]
@@
[…]
$2.\displaystyle=\penalty 10000\ M(A)+M(B)-k(2k)\penalty 10000\ \geq\penalty 100002k^{2}+2k^{2}-2k^{2}\penalty 10000\ =\penalty 100002k^{2}\,$.
@@
Proof. We present a construction for (i), which we will modify for (ii) afterwards.
@@
Let $a,b,c\in\mathbb{N}$ which differ by at most $1$ with a+b+c=k1a+b+c=k-1. Let $V_{0}$ be an anticlique on $k$ vertices and let $V_{1},\dots,V_{l}$ be disjoint cliques on k1k-1 vertices that are also disjoint to $V_{0}$, where $l=3+2a+b$. We partition each $V_{j}$ for $j\geq 1$ into three cliques $V_{j}=A_{j}\cup B_{j}\cup C_{j}$ with $|A_{j}|=a$, $|B_{j}|=b$ and $|C_{j}|=c$.
Let $w_{4},w_{5},\dots,w_{l}$ be a list of the l3=2a+bl-3=2a+b vertices in $W:=A_{1}\cup B_{1}\cup A_{2}$. For $j\in\{1,\dots,l\}$, we add all possible edges between $V_{j}$ and some set $W_{j}\subseteq V_{0}\cup\dots\cup V_{j-1}$ of $k$ vertices. We set $W_{j}=V_{0}$ for $j\in\{1,2,3\}$ and $W_{j}=A_{j-1}\,\cup\,B_{j-2}\,\cup\,C_{j-3}\,\cup\,\{w_{j}\}$ otherwise. For $j\in\{1,2,3,l\}$, this is shown in Figure 1.
@@
$|X|\penalty 10000\ =\penalty 10000\ c+b+c+(k-1)\penalty 10000\ \leq\penalty 10000\ 1+a+b+c+(k-1)\penalty 10000\ =\penalty 10000\ 2k-1\,.$
Figure 1: The graph $H$. Only the edges between $V_{j}$ and $W_{j}$ for $j\in\{1,2,3,l\}$ are shown.
@@
Now let $j\geq 4$ and assume that there is a $(k+1)$-connected subgraph $G^{\prime}$ on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices in $G_{j}$. Since the neighborhood $W_{j}$ of $V_{j}$ in $G_{j}$ consists of only $k$ vertices, $G^{\prime}$ is either a subgraph of $G_{j}-V_{j}=G_{j-1}$ or of $G_{j}[V_{j}\cup W_{j}]$. The first case is prevented by the induction hypothesis, so assume the second. Since there are no edges between each two of the sets $A_{j-1}$, $B_{j-2}$ and $C_{j-3}$, and those three are separated after deleting the set $V_{j}\cup\{w_{j}\}$ of $k$ vertices, $V(G^{\prime})$ must be completely contained in one of the sets $V_{j}\cup\{w_{j}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j}\}\cup C_{j-3}$. But all of them have size at most $(k-1)+1+\big\lceil\frac{k-1}{3}\big\rceil\leq\big\lceil\frac{4}{3}k\big\rceil$, a contradiction.
This proves the claimed properties of $H$. We now inductively construct a series $H_{0},H_{1},\dots$ of graphs without a $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices such that in $H_{j}$ there is a set $X_{j}$ of $\max\{2k-2^{j},0\}$ vertices of degree at least 2k22k-2 and all vertices in $V(H_{j})\setminus X_{j}$ have degree at least 3k33k-3. Then for $j\geq\log_{2}2k$, $H_{j}$ has minimum degree at least 3k33k-3.
Let $H_{0}=H$ and $X_{0}=X$ (possibly replenished to size 2k12k-1), which satisfy the conditions. Now assume we have already constructed $H_{j},X_{j}$ and we want to construct $H_{j+1},X_{j+1}$. Let $Y_{j}\subseteq X_{j}$ with $|Y_{j}|=\min\{k,|X_{j}|\}$. Let $H^{\prime}_{j}$ be a copy of $H_{j}$. We define $H_{j+1}$ as the union of $H_{j}$ and $H^{\prime}_{j}$ when we identify $Y_{j}$ and its copy in $H^{\prime}_{j}$, that is, the subgraphs $H_{j}$ and $H^{\prime}_{j}$ of $H_{j+1}$ intersect only in the vertex set $Y_{j}$. Since $|Y_{i}|\leq k$ and both $H_{j}$ and $H^{\prime}_{j}$ contain no $(k+1)$-connected subgraph on more than $\big\lceil\frac{4}{3}k\big\rceil$ vertices, neither $H_{j+1}$ does. For every $y\in Y_{j}$, we have
$d_{H_{j+1}}(y)\penalty 10000\ =\penalty 10000\ d_{H_{j}}(y)+d_{H^{\prime}_{j}}(y)-d_{H_{j}[Y_{j}]}(y)\penalty 10000\ \geq\penalty 10000\ 2(k-1)+2(k-1)-(k-1)\penalty 10000\ =\penalty 10000\ 3(k-1)\,.$
Also the vertices that are not contained in $X_{j}$ or its copy in $H_{j}^{\prime}$ have degree at least 3k33k-3. The remaining vertices from $X_{j}\setminus Y_{j}$ and its copy in $H_{j}^{\prime}$ form the new set $X_{j+1}$ of size
[…]
@@
For (ii), we replace the complete graphs on each of the sets $A_{j},B_{j},C_{j}$ for $j\geq 1$ by the disjoint union of two complete graphs of size as equal as possible. By doing so, we loose in $H_{0}=H$ a degree $\frac{1}{6}k+\mathcal{O}(1)$ at every vertex. One observes inductively for $j=0,1,\dots$ that $d_{H_{j}[Y_{j}]}(v)\leq d_{H_{j}[X_{j}]}(v)\leq k-\frac{1}{6}k+\mathcal{O}(1)$ for every vertex $v\in Y_{j}$. Hence indeed
[…]
The proof of the non-existence of a $(k+1)$-connected subgraph first works as for (i): The vertex set of a $(k+1)$-connected subgraph $G^{\prime}$ must be completely contained in one of the sets $V_{j}\cup\{w_{j}\}\cup A_{j-1}$,  $V_{j}\cup\{w_{j}\}\cup B_{j-2}$  or $V_{j}\cup\{w_{j}\}\cup C_{j-3}$, for some $j\geq 4$. Assume $V(G^{\prime})\subseteq V_{j}\cup\{w_{j}\}\cup A_{j-1}$. The two components of $H[A_{j-1}]$ can be separated by deleting the set $V_{j}\cup\{w_{j}\}$ of $k$ vertices. Hence $G^{\prime}$ cannot contain vertices from both components. The same holds for $A_{j}\subseteq V_{j}$. Together, $V(G^{\prime})\leq 2\big\lceil\frac{a}{2}\big\rceil+b+c+1\leq 1+a+b+c+1=k+1$, which is too low for being $(k+1)$-connected. The other two cases work similarly. $\Box$
@@
There is a $C<3$ such that, for all $k\in\mathbb{N}$, every graph of minimum degree at least Ck1Ck-1 has a $(k+1)$-connected subgraph.
By Theorem On highly connected subgraphs, a suitable constant must satisfy $C\geq 3-\frac{1}{6}$. The conjecture is posed in a way that allows to assume that $k\geq k_{0}$ for some sufficiently large $k_{0}$: For $k<k_{0}$, $\big\lceil\big(3-\frac{1}{k_{0}}\big)k-1\big\rceil\,=\,\big\lceil 3k-\frac{k}{k_{0}}-1\big\rceil\,=\,3k-1$, so Theorem On highly connected subgraphs suffices for $C=3-\frac{1}{k_{0}}<3$.
@@
Let $k\geq 1$. Every triangle-free graph with average degree at least $2k$ has a $(k+1)$-connected subgraph (on at least $2(k+1)$ vertices).
Proof. First, we prove the well-known statement that, if $G$ is a triangle-free graph on $2l$ vertices, then $e(G)\leq l^{2}$. This is true for $l=1$. If $l>1$ and $E(G)\neq\emptyset$, consider an edge $vw\in E(G)$. Since $G$ is triangle-free, the neighborhoods of $v$ and $w$ are disjoint and hence $d_{G}(v)+d_{G}(w)\leq v(G)=2l$. This yields $e(G)=e(G-v-w)+d_{G}(v)+d_{G}(w)-1\linebreak\leq(l-1)^{2}+2l-1=l^{2}$.
@@
Now assume $v(G)>2k$. $G$ itself is not $(k+1)$-connected. Hence it has a $k$-separation $(A,B)$ (see Theorem On highly connected subgraphs). By symmetry, we may assume that $v(A)\geq v(B)$.
@@
$e(G)\penalty 10000\ =\penalty 10000\ e(G-b)+d_{G}(b)\penalty 10000\ \leq\penalty 10000\ e(G-b)+k\penalty 10000\ \leq\penalty 10000\ k\big(v(G)-1-k\big)+k\penalty 10000\ =\penalty 10000\ k\big(v(G)-k\big)\,.$
@@
$e(G)\penalty 10000\ \leq\penalty 10000\ e(A)+e(B)\penalty 10000\ \leq\penalty 10000\ k\big(v(A)-k\big)+k\big(v(B)-k\big)\penalty 10000\ =\penalty 10000\ k\big(v(G)-k\big)\,.$
@@
Already in [3], Mader proved the above result for graphs of girth larger than $k+1$. He also presented a construction of bipartite graphs of arbitrarily large girth and average degree arbitrarily close to $2k$ without a $(k+1)$-connected subgraph. Even the minimum degree version of Theorem On highly connected subgraphs is best possible, that is, there exists a bipartite graph with minimum degree 2k12k-1 without a $(k+1)$-connected subgraph. One can build such a graph by gluing together copies of $H_{0}:=K_{k,2k-1}$ similarly to the final step of the proof of Theorem On highly connected subgraphs.
@@
Mader proved in [4] that every (simple) graph $G$ with $e(G)>k\big(v(G)-\frac{k+1}{2}\big)$ and $v(G)\geq k$ contains a $(k+1)$-edge-connected subgraph. This condition is sharp for every $k$ and every $v(G)$: Consider the graph $G$ on a clique $A$ on $k$ vertices and an anticlique on $V(G)\setminus A$ with all possible edges between $A$ and $V(G)\setminus A$. Clearly, $G$ has not even a subgraph of minimum degree at least $k+1$. Note that, when $v(G)$ tends to $\infty$, the average degree of $G$ tends to $2k$.
We first show that an average degree of at least $2k$ has some stronger implications. For (simple) graphs in particular, it implies the existence of a $(k+1)$-edge-connected subgraph on more than $2k$ vertices.
@@
Let $G$ be a loopless multigraph with e(G)>k(v(G)1)e(G)\,>\,k\,(v(G)-1). Then $G$ contains a submultigraph $G^{\prime}$ with
@@
    If $G$ is a (simple) graph, then $v(G^{\prime})>2k$.
Proof. Let $G^{\prime}$ be a smallest nonempty induced submultigraph of $G$ with $e(G^{\prime})\,>\,k\,(v(G^{\prime})-1)$. Let $\mathcal{P}$ be a partition of $V(G^{\prime})$ in more than one class. Then for all $A\in\mathcal{P}$, $e(G[A])\leq k(|A|-1)$. Hence the number of edges between distinct classes of $\mathcal{P}$ satisfies
[…]
$=\penalty 10000\ k\Big(v(G^{\prime})-1+|\mathcal{P}|-\sum_{A\in\mathcal{P}}|A|\Big)\penalty 10000\ =\penalty 10000\ k(|\mathcal{P}|-1)\,.$
In particular, it follows from the case $|\mathcal{P}|=2$, that $G^{\prime}$ is $(k+1)$-edge-connected. Moreover, by the Theorem by Tutte [5] and Nash-Williams [6], $G^{\prime}$ contains $k$ pairwise edge-disjoint spanning trees. If $G$ and hence $G^{\prime}$ is a (simple) graph, then $G^{\prime}$ has more than $2k$ vertices, since otherwise $e(G^{\prime})\,\leq\,\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)\,\leq\,k\,(v(G^{\prime})-1)$. $\Box$
@@
    Every (2k1)(2k-1)-regular (simple) graph $G$ without a complete graph on $2k$ vertices as a component does not contain a subgraph $G^{\prime}$ with $k$ pairwise edge-disjoint spanning trees, because every subgraph $G^{\prime}$ has less than the necessary $k(v(G^{\prime})-1)$ edges:
@@
    • If $v(G^{\prime})=2k$, then $G^{\prime}$ is not complete, since otherwise it would be a component of $G$. Hence $e(G^{\prime})<\frac{1}{2}v(G^{\prime})(v(G^{\prime})-1)=k(v(G^{\prime})-1)$.